# Latent Adversarial Regularization for Offline Preference Optimization

### v1

@@ Abstract
Learning from human feedback typically relies on preference optimization that constrains policy updates through token-level regularization. However, preference optimization for language models is particularly challenging because token-space similarity does not imply semantic or behavioral similarity. To address this challenge, we leverage latent-space regularization for language model preference optimization. We introduce GANPO, which achieves latent-space regularization by penalizing divergence between the internal representations of a policy model and a reference model. Given that latent representations are not associated with explicit probability densities, we adopt an adversarial approach inspired by GANs to minimize latent-space divergence. We integrate GANPO as a regularizer into existing offline preference optimization objectives. Experiments across multiple model architectures and tasks show consistent improvements from latent-space regularization. Further, by comparing GANPO-induced inferential biases with those from token-level regularization, we find that GANPO provides more robust structural feedback under distributional shift and noise while maintaining comparable downstream performance with minor computational overhead.
@@ 3.2 Generative Adversarial Formulation
$\mathbb{D}_{\mathrm{Ra}}(p_{\theta}\,\|\,p_{\mathrm{ref}})$ (12)
$=\sup_{\phi}\Big[\mathbb{E}_{h_{\mathrm{ref}}}\big[\log\tilde{D}_{\phi}(h_{\mathrm{ref}})\big]+\mathbb{E}_{h_{\theta}}\big[\log\big(1-\tilde{D}_{\phi}(h_{\theta})\big)\big]\Big].$ (13)
@@
$\mathbb{D}_{\mathrm{Ra}}(p_{\theta}\,\|\,p_{\mathrm{ref}})$ $=\sup_{\phi}\ -\mathrm{BCE}_{\phi}(h_{\mathrm{ref}},h_{\theta})$ (16)
$=-\inf_{\phi}\ \mathrm{BCE}_{\phi}(h_{\mathrm{ref}},h_{\theta}).$ (17)
@@ 5.2 Results and Analysis
(2) IFEval (zhou2023instructionfollowing): To assess the stability of instruction following under noise, we measure the strict prompt-level accuracy with structured outputs on IFEval (Figure 3(b)). Here, the contrast is stark: DPO suffers from more severe structural collapse as temperature increases (dropping nearly 20% in accuracy from $T=0.5$ to $T=1.0$), indicating that its adherence to constraints heavily relies on greedy decoding. In contrast, GANPO demonstrates stronger resilience, retaining good strict accuracy under stochastic sampling.
@@ Appendix A Relativistic Average Divergence
$\mathbb{D}_{\mathrm{Ra}}(p_{\theta}\,\|\,p_{\mathrm{ref}}):=\sup_{\phi}\Big[\mathbb{E}_{h_{\mathrm{ref}}}\big[\log\sigma\big(C_{\phi}(h)-m_{\theta}\big)\big]+\mathbb{E}_{h_{\theta}}\big[\log\big(1-\sigma\big(C_{\phi}(h)-m_{\mathrm{ref}}\big)\big)\big]\Big],$ (22)
@@
$\mathbb{D}_{\mathrm{Ra}}(p_{\theta}\,\|\,p_{\mathrm{ref}}):=\sup_{C:\mathcal{H}\to\mathbb{R}}\Big[\mathbb{E}_{h_{\mathrm{ref}}}\big[\log\sigma\big(C(h)-m_{\theta}\big)\big]+\mathbb{E}_{h_{\theta}}\big[\log\big(1-\sigma\big(C(h)-m_{\mathrm{ref}}\big)\big)\big]\Big].$ (24)
@@
$\mathbb{D}_{\mathrm{Ra}}(p_{\theta}\,\|\,p_{\mathrm{ref}})$ $:=\sup_{C:\mathcal{H}\to\mathbb{R}}\Big[\mathbb{E}_{h_{\mathrm{ref}}}\big[\log\sigma\big(C(h)-m_{\theta}\big)\big]+\mathbb{E}_{h_{\theta}}\big[\log\big(1-\sigma\big(C(h)-m_{\mathrm{ref}}\big)\big)\big]\Big].$ (26)
$=\sup_{C:\mathcal{H}\to\mathbb{R}}\Big[\mathbb{E}_{h_{\mathrm{ref}}}\big[\log\sigma\big(C(h)-m_{\theta}\big)\big]+\mathbb{E}_{h_{\theta}}\big[\log\sigma\big(m_{\mathrm{ref}}-C(h)\big)\big)\big]\Big]$ (27)

### v2

@@ Abstract
Learning from human feedback typically relies on preference optimization that constrains policy updates through token-level regularization. However, preference optimization for language models is particularly challenging because token-space similarity does not imply semantic or behavioral similarity. To address this challenge, we leverage latent-space regularization for language model preference optimization. We introduce GANPO, which achieves latent-space regularization by penalizing divergence between the internal representations of a policy model and a reference model. Given that latent representations are not associated with explicit probability densities, we adopt an adversarial approach inspired by GANs to minimize latent-space divergence. We integrate GANPO as a regularizer into existing offline preference optimization objectives. Experiments across multiple model architectures and tasks show consistent improvements from latent-space regularization. Further, by comparing GANPO-induced inferential biases with those from token-level regularization, we find that GANPO provides more robust structural feedback under distributional shift and noise while maintaining comparable downstream performance with minor computational overhead. Our code is available at https://github.com/enyijiang/GANPO.
@@ 3.2 Generative Adversarial Formulation
$\mathbb{D}_{\mathrm{Ra}}(p_{\theta}\,\|\,p_{\mathrm{ref}})=\log 4$ (12)
$+\sup_{\phi}\Big[\mathbb{E}_{h_{\mathrm{ref}}}\big[\log\tilde{D}_{\phi}(h_{\mathrm{ref}})\big]+\mathbb{E}_{h_{\theta}}\big[\log\big(1-\tilde{D}_{\phi}(h_{\theta})\big)\big]\Big].$ (13)
@@
$\mathbb{D}_{\mathrm{Ra}}(p_{\theta}\,\|\,p_{\mathrm{ref}})$ $=\log 4+\sup_{\phi}\ -\mathrm{BCE}_{\phi}(h_{\mathrm{ref}},h_{\theta})$ (16)
$=\log 4-\inf_{\phi}\ \mathrm{BCE}_{\phi}(h_{\mathrm{ref}},h_{\theta}).$ (17)
@@ 5.2 Results and Analysis
(2) IFEval (zhou2023instructionfollowing): To assess the stability of instruction following under noise, we measure the strict prompt-level accuracy with structured outputs on IFEval (Figure 3(b)). Here, the contrast is stark: DPO suffers from more severe structural collapse as temperature increases (dropping nearly 20% in accuracy from $T=0.0$ to $T=1.0$), indicating that its adherence to constraints heavily relies on greedy decoding. In contrast, GANPO demonstrates stronger resilience, retaining good strict accuracy under stochastic sampling.
@@ 8 Impact Statement
9 Acknowledgments
SK acknowledges support by NSF 2046795 and 2205329, IES R305C240046, ARPA-H, the MacArthur Foundation, Schmidt Sciences, HAI, OpenAI, Microsoft, and Google. This research used the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois. DeltaAI is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications. Further, we gratefully acknowledge the use of GPU computing resources provided by the CAIS Compute Cluster at the Center for AI Safety (Safe.ai).
@@ Appendix A Relativistic Average Divergence
$\mathbb{D}_{\mathrm{Ra}}(p_{\theta}\,\|\,p_{\mathrm{ref}}):=\sup_{\phi}\Big[\mathbb{E}_{h_{\mathrm{ref}}}\big[\log\sigma\big(C_{\phi}(h)-m_{\theta}\big)\big]+\mathbb{E}_{h_{\theta}}\big[\log\big(1-\sigma\big(C_{\phi}(h)-m_{\mathrm{ref}}\big)\big)\big]\Big]+\log 4,$ (22)
@@
$\mathbb{D}_{\mathrm{Ra}}(p_{\theta}\,\|\,p_{\mathrm{ref}}):=\sup_{C:\mathcal{H}\to\mathbb{R}}\Big[\mathbb{E}_{h_{\mathrm{ref}}}\big[\log\sigma\big(C(h)-m_{\theta}\big)\big]+\mathbb{E}_{h_{\theta}}\big[\log\big(1-\sigma\big(C(h)-m_{\mathrm{ref}}\big)\big)\big]\Big]+\log 4.$ (24)
@@
$\mathbb{D}_{\mathrm{Ra}}(p_{\theta}\,\|\,p_{\mathrm{ref}})$ $:=\sup_{C:\mathcal{H}\to\mathbb{R}}\Big[\mathbb{E}_{h_{\mathrm{ref}}}\big[\log\sigma\big(C(h)-m_{\theta}\big)\big]+\mathbb{E}_{h_{\theta}}\big[\log\big(1-\sigma\big(C(h)-m_{\mathrm{ref}}\big)\big)\big]\Big]+\log 4.$ (26)
$=\sup_{C:\mathcal{H}\to\mathbb{R}}\Big[\mathbb{E}_{h_{\mathrm{ref}}}\big[\log\sigma\big(C(h)-m_{\theta}\big)\big]+\log 2+\mathbb{E}_{h_{\theta}}\big[\log\sigma\big(m_{\mathrm{ref}}-C(h)\big)\big)\big]+\log 2\Big]$ (27)